# EmT: A Novel Transformer for Generalized Cross-subject EEG Emotion Recognition

### v1

@@ Abstract
Integrating prior knowledge of neurophysiology into neural network architecture enhances the performance of emotion decoding. While numerous techniques emphasize learning spatial and short-term temporal patterns, there has been limited emphasis on capturing the vital long-term contextual information associated with emotional cognitive processes. In order to address this discrepancy, we introduce a novel transformer model called emotion transformer (EmT). EmT is designed to excel in both generalized cross-subject EEG emotion classification and regression tasks. In EmT, EEG signals are transformed into a temporal graph format, creating a sequence of EEG feature graphs using a temporal graph construction module (TGC). A novel residual multi-view pyramid GCN module (RMPG) is then proposed to learn dynamic graph representations for each EEG feature graph within the series, and the learned representations of each graph are fused into one token. Furthermore, we design a temporal contextual transformer module (TCT) with two types of token mixers to learn the temporal contextual information. Finally, the task-specific output module (TSO) generates the desired outputs. Experiments on four publicly available datasets show that EmT achieves higher results than the baseline methods for both EEG emotion classification and regression tasks. The code is available at https://github.com/yi-ding-cs/EmT
@@ I Introduction
Emotion recognition using electroencephalography (EEG) plays an important role in brain-computer interface (BCI) assisted mental disorder regulation. It requires machine to perceive human emotional states from brain activities using artificial intelligence techniques [1, 2]. The predictions of emotions can be used for neurofeedback in the regulation process [3]. Accurate predictions and fair generalization abilities to unseen subjects are crucial in building robust real-world BCI systems. Deep learning methods have shown promising results for accurate detection of brain activities [4, 5]. The design of the neural network architecture becomes crucial.
Incorporating neurophysiological prior knowledge into neural network architectures improves emotion decoding performance [6, 7, 8]. Common considerations include left-right hemisphere asymmetries, neurophysiologically meaningful graph connections, and the temporal dynamics of EEG signals. According to neuropsychological studies, the left and right hemispheres react differently to emotions, particularly in the frontal areas [9]. Several deep learning methods [8, 10, 11] draw inspiration from this prior knowledge, achieving improved emotion decoding performance. EEG electrodes are placed on the scalp, which naturally forms a non-Euclidean structure. Therefore, many studies treat EEG signals as graphs, using either a neurophysiologically designed adjacency matrix [7] or a learnable one [12, 13]. However, there is still more prior knowledge that should be investigated. As one of the high-order cognitive processes in the brain, emotion consists of more basic processes such as attentional, perceptual, and mnemonic system processes [14]. Different brain regions are cooperatively activated under different cognitive processes, e.g., frontal and parietal networks in attention [15] and medial temporal lobes, prefrontal cortex, and parietal cortex interactions for episodic memory [16]. Using a pre-defined or a single learnable adjacency matrix cannot capture complex brain region connectives underlying emotional processes. Another prior knowledge about emotions is that emotional states are continuous in short periods while not consistent along the long stimuli [17]. Less attention is paid to this temporal contextual information underlying emotions.
To address the above-mentioned problems, we propose a novel transformer-based structure, named emotion transformer (EmT), for both generalized cross-subject emotion classification and regression tasks. To the best of our knowledge, this is the first work to explore transformer-based structures on cross-subject EEG emotion classification and regression tasks together. To learn temporal contextual information, we represent EEG segments as temporal graphs, as shown in Figure. 2. We learn the spatial information of each EEG graph to form a token. Then the long-short time contextual information is learned upon the token sequence.
@@
We further propose temporal contextual transformers (TCT) with different token mixers [19] to learn contextual information from the token sequence for EEG emotion classification and regression tasks. An EEG trial refers to the period when one type of stimuli, e.g., a happy movie clip, is presented to the subject while the EEG is recorded. The label of the entire trial is assigned as happy according to either the self-assessments or the stimulus contents in classification tasks. However, the emotional states of the subjects change along with the stimuli [17]. Cutting each trial into short segments and assigning the same label to them will induce noisy labels [7]. To relieve noisy label issues and learn temporal contextual information of emotions, we can use a longer sliding window with short-shifting steps to split each trial, and the longer segments are further cut into sub-segments. Multi-head self-attention (MSA) in transformers [20] can attentively emphasis the parts that are highly correlated to the overall emotional state of the longer EEG segment. Because emotion is continuous [21], the underlying state is consistent in a short period. Hence, we propose a short-time aggregation (STA) layer after MSA to learn the long-short-time contextual information. Different from classification tasks, the label is temporally continuous in regression tasks. The model needs to regress the continuous changes of the emotional states for all the segments within a sequence. Although MSA can globally emphasis important parts in the sequence, recurrent neural networks (RNN) can further fuse the information of all the segments recurrently. This global information fusion ability makes RNN more suitable for regression task. Hence, we propose to use an RNN-based token mixer in TCT for the regression tasks instead of MSA. To this end, we propose TCT-Clas and TCT-Regr as the token mixers in EmT for EEG emotion classification and regression tasks.
@@
  • We propose a residual multi-view pyramid GCN module (RMPG) to learn multi-view and multi-level graph embeddings incorporating the neuroscience knowledge that emotion consists of multiple basic cognitive processes.
  • Two types of temporal contextual transformer (TCT) blocks with task-specific token mixers are proposed to learn temporal contextual information from EEG.
@@ III Method
In this work, we propose a novel transformer, EmT, for both the generalized cross-subject EEG emotion classification and regression tasks. The network architecture is shown in Figure  1. EmT consists of four main parts: (1) temporal graph construction module (TGC), (2) RMPG, (3) TCT, and (4) task-specific output module (TSO). In TGC, EEG signals are transformed into temporal graph format that is a sequence of EEG feature graphs, as shown in Figure  2. RMPG learns dynamical graph representations for each EEG feature graph within the series and the learned representations of each graph are fused into one token. TCT learns the temporal contextual information via specially designed token mixers. Finally, the TSO module will generate the desired output for classification and regression tasks accordingly.
@@ III-A EEG-temporal-graph Representations
Temporal graphs are constructed to allow the neural network to learn spatial and temporal contextual information. Two steps are required to generate the temporal graphs from EEG, which are EEG segment/sub-segment segmentation and feature extraction, respectively. Figure  2 shows the construction process for one EEG segment.
@@ III-B Residual Multi-view Pyramid GCN
To modulate different brain region connections for multiple basic cognitive processes underlying emotions [14], we propose to use multiple different-layer GCNs, $\Phi_{g}^{0}(\cdot),\Phi_{g}^{1}(\cdot),...,\Phi_{g}^{i}(\cdot)$} with learnable adjacency matrices [12], $\bm{A}^{0},\bm{A}^{1},...,\bm{A}^{i}$}. Each $\bm{A}^{i}$ can learn one view of graph connections that belongs to a certain basic cognitive process. For each $\Phi_{g}^{i}(\cdot)$, stacking different layers of GCNs can learn different degrees of node cluster similarity [34]. Intuitively, for localized connections, such as electrodes within a brain functional area, a deeper GCN can get a consistent representation among these nodes. However, for global connections that are among different brain functional areas, a shallow GCN can aggregate the information among these areas while not over smooth it. The deeper the $\Phi_{g}^{i}(\cdot)$ is, the taller the feature pyramid level its output has. A linear projection layer, $\textrm{LP}(\cdot)$, is added for each $\Phi_{g}^{i}(\cdot)$ to project the flattened graph representations into a hidden embedding denoted by $\bm{H}_{g}^{i}\in\mathbb{R}^{d_{g}}$. By stacking $\bm{H}_{g}^{i}$ from parallel different-layer GCNs, we can get multi-view pyramid graph embeddings as:
@@ III-C1 Token Mixers for Classification Tasks
Considering the fact that emotion is short-term continuous and long-term varying [17], we propose a short-time aggregation (STA) layer after MSA to learn the long-short-time contextual information. A dropout layer with a scaling factor $\alpha$, $\textrm{dp}(\alpha)$ is applied to control the forgetting rate of the temporal context. $\alpha$ is a hyper-parameter that will scale down the overall dropout rate in STA. To capture short-term consist patterns, CNN kernels denoted by $\bm{K}_{cnn}$ whose size and step are $(n_{anchor},1)$, $(1,1)$ are utilized to aggregate $n_{anchor}$ temporal neighbors after MSA. Let $\bm{H}_{attn}\in\mathbb{R}^{n_{head}\times seq\times d_{head}}$ denotes one output of MSA. STA can be described as:
@@ V-A Emotion Classification
The observations on the THU-EP and FACED datasets are different. For the THU-EP dataset, EmT-B achieves the best F1 score of 0.724, while Conformer achieves the best accuracy (0.601). On the FACED dataset, EmT-B leads with the second-best accuracy (0.608) and the highest F1 score (0.740). AMDET achieves the second-best F1 score of 0.726, while the best accuracy is achieved by TSception (0.619). As the classes are imbalanced in both the THU-EP and FACED datasets, F1 scores are more important than accuracy. EmT-B achieves the highest F1 scores on both THU-EP and FACED, demonstrating its effectiveness in emotion classification. Different from the observations on the SEED dataset, the baselines using EEG as input achieve better performance than those using features as input. This might be because there are more subjects in THU-EP and FACED, and using EEG directly can provide more information. Overall, the EmT series methods demonstrate the best overall performance across the evaluated datasets.
@@ V-D Effect of EEG Features
Figure 3 illustrates the impact of different feature types on the SEED dataset for classification tasks and the MAHNOB-HCI dataset for regression tasks. We compared three types of features: PSD, DE, and rPSD, evaluating their effects on accuracy and F1 score for classification tasks, as well as on MRSE, CCC, and PCC for regression tasks. For classification tasks, rPSD outperformed the other two feature types, providing the highest accuracy and F1 score. Specifically, rPSD achieved 5.9% and 11.5% higher accuracy and F1 score, respectively, compared to EmT using DE. Furthermore, rPSD showed a 13.0% and 18.9% improvement in accuracy and F1 score, respectively, over PSD. These results suggest that rPSD is a superior spectral feature for EEG emotion classification tasks. For regression tasks, using rPSD resulted in better MRSE and CCC compared to using DE. However, the difference in PCC between rPSD and DE was minimal, with rPSD achieving a PCC of 0.491 and DE achieving a PCC of 0.490. Notably, when PSD was used as a feature, the model failed to converge, so we excluded these results from further analysis. Based on these findings, rPSD is demonstrated to be a more effective feature than DE, and both rPSD and DE are superior to PSD for EEG emotion recognition tasks.
@@
V-E Effect of The Number of TCT Blocks
We vary the number of TCT blocks from 2 to 8 and monitor their effect on the results of the two tasks. The results are shown in Figure 4. For classification, shown in Figure 4 (a), as the number of TCT blocks increases from 2 to 8, there is a noticeable improvement in both ACC and F1 scores. Specifically, accuracy increases from 0.780 to 0.802, while the F1 score shows a more substantial rise from 0.759 to 0.821, with the most significant improvement occurring between 6 and 8 TCT blocks. This indicates that adding more TCT blocks enhances the model’s ability to capture temporal contextual information, thereby improving classification performance. Conversely, for regression tasks, shown in Figure 4 (b), the number of TCT blocks has little to no effect on the performance metrics. The RMSE remains stable around 0.06, and both the PCC and CCC show minimal variation, hovering around 0.48 and 0.39, respectively. This suggests that while increasing TCT blocks benefits classification by improving the capture of temporal contextual information, it does not significantly impact regression performance.

### v2

@@ Abstract
Integrating prior knowledge of neurophysiology into neural network architecture enhances the performance of emotion decoding. While numerous techniques emphasize learning spatial and short-term temporal patterns, there has been limited emphasis on capturing the vital long-term contextual information associated with emotional cognitive processes. In order to address this discrepancy, we introduce a novel transformer model called emotion transformer (EmT). EmT is designed to excel in both generalized cross-subject EEG emotion classification and regression tasks. In EmT, EEG signals are transformed into a temporal graph format, creating a sequence of EEG feature graphs using a temporal graph construction module (TGC). A novel residual multi-view pyramid GCN module (RMPG) is then proposed to learn dynamic graph representations for each EEG feature graph within the series, and the learned representations of each graph are fused into one token. Furthermore, we design a temporal contextual transformer module (TCT) with two types of token mixers to learn the temporal contextual information. Finally, the task-specific output module (TSO) generates the desired outputs. Experiments on four publicly available datasets show that EmT achieves higher results than the baseline methods for both EEG emotion classification and regression tasks. The code is available at https://github.com/yi-ding-cs/EmT.
@@ I Introduction
Emotion recognition using electroencephalography (EEG) plays an important role in brain-computer interface (BCI) assisted mental disorder regulation. It requires machine to perceive human emotional states from brain activities using artificial intelligence techniques [1, 2]. The predictions of emotions can be used for neurofeedback in the regulation process [3]. Accurate predictions and fair generalization abilities to unseen subjects are crucial in building robust real-world BCI systems. Deep learning methods have shown promising results for accurate detection of brain activities [4, 5, 10763464]. The design of the neural network architecture becomes crucial.
Incorporating neurophysiological prior knowledge into neural network architectures improves emotion decoding performance [6, 7, 8, 10506986]. Common considerations include left-right hemisphere asymmetries, neurophysiologically meaningful graph connections, and the temporal dynamics of EEG signals. According to neuropsychological studies, the left and right hemispheres react differently to emotions, particularly in the frontal areas [9]. Several deep learning methods [8, 10, 11] draw inspiration from this prior knowledge, achieving improved emotion decoding performance. EEG electrodes are placed on the scalp, which naturally forms a non-Euclidean structure. Therefore, many studies treat EEG signals as graphs, using either a neurophysiologically designed adjacency matrix [7] or a learnable one [12, 13, 10340644]. However, there is still more prior knowledge that should be investigated. As one of the high-order cognitive processes in the brain, emotion consists of more basic processes such as attentional, perceptual, and mnemonic system processes [14]. Different brain regions are cooperatively activated under different cognitive processes, e.g., frontal and parietal networks in attention [15] and medial temporal lobes, prefrontal cortex, and parietal cortex interactions for episodic memory [16]. Using a pre-defined or a single learnable adjacency matrix cannot capture complex brain region connectives underlying emotional processes. Another prior knowledge about emotions is that emotional states are continuous in short periods while not consistent along the long stimuli [17]. Less attention is paid to this temporal contextual information underlying emotions.
To address the above-mentioned problems, we propose a novel transformer-based structure, named emotion transformer (EmT), for both generalized cross-subject emotion classification and regression tasks. To the best of our knowledge, this is the first work to explore transformer-based structures on cross-subject EEG emotion classification and regression tasks together. To learn temporal contextual information, we represent EEG segments as temporal graphs, as shown in Figure 2. We learn the spatial information of each EEG graph to form a token. Then the long-short time contextual information is learned upon the token sequence.
@@
We further propose temporal contextual transformers (TCT) with different token mixers [19] to learn contextual information from the token sequence for EEG emotion classification and regression tasks. An EEG trial refers to the period when one type of stimuli, e.g., a happy movie clip, is presented to the subject while the EEG is recorded. The label of the entire trial is assigned as happy according to either the self-assessments or the stimulus contents in classification tasks. However, the emotional states of the subjects change along with the stimuli [17]. Cutting each trial into short segments and assigning the same label to them will induce noisy labels [7]. To relieve noisy label issues and learn temporal contextual information of emotions, we can use a longer sliding window with short-shifting steps to split each trial, and the longer segments are further cut into sub-segments. Multi-head self-attention (MSA) in transformers [20] can attentively emphasis the parts that are highly correlated to the overall emotional state of the longer EEG segment. Because emotion is continuous [21], the underlying state is consistent in a short period. Hence, we propose a short-time aggregation (STA) layer after MSA to learn the long-short-time contextual information. Different from classification tasks, the label is temporally continuous in regression tasks. The model needs to regress the continuous changes of the emotional states for all the segments within a sequence. Although MSA can globally emphasis important parts in the sequence, recurrent neural networks (RNN) can further fuse the information of all the segments recurrently. This causal information fusion ability makes RNN more suitable for regression task. Hence, we propose to use an RNN-based token mixer in TCT for the regression tasks instead of MSA. To this end, we propose TCT-Clas and TCT-Regr as the token mixers in EmT for EEG emotion classification and regression tasks.
@@
  • We introduce a residual multi-view pyramid GCN (RMPG) module, designed to learn multi-view and multi-level graph embeddings using multiple learnable adjacency matrices. This approach incorporates neuroscience knowledge, recognizing that emotions are composed of various basic cognitive processes.
  • Two types of temporal contextual transformer (TCT) blocks with task-specific token mixers are proposed to capture temporal contextual information from EEG data, accounting for the differing learning objectives between classification and regression tasks.
@@ III Method
In this work, we propose a novel transformer, EmT, for both the generalized cross-subject EEG emotion classification and regression tasks. The network architecture is shown in Figure 1. EmT consists of four main parts: (1) temporal graph construction module (TGC), (2) RMPG, (3) TCT, and (4) task-specific output module (TSO). In TGC, EEG signals are transformed into temporal graph format that is a sequence of EEG feature graphs, as shown in Figure 2. RMPG learns dynamical graph representations for each EEG feature graph within the series and the learned representations of each graph are fused into one token. TCT learns the temporal contextual information via specially designed token mixers. Finally, the TSO module will generate the desired output for classification and regression tasks accordingly.
@@ III-A EEG-temporal-graph Representations
Temporal graphs are constructed to allow the neural network to learn spatial and temporal contextual information. Two steps are required to generate the temporal graphs from EEG, which are EEG segment/sub-segment segmentation and feature extraction, respectively. Figure 2 shows the construction process for one EEG segment.
@@ III-B Residual Multi-view Pyramid GCN
To modulate different brain region connections for multiple basic cognitive processes underlying emotions [14], we propose to use multiple different-layer GCNs, $\Phi_{g}^{0}(\cdot),\Phi_{g}^{1}(\cdot),...,\Phi_{g}^{i}(\cdot)$} with learnable adjacency matrices [12], $\bm{A}^{0},\bm{A}^{1},...,\bm{A}^{i}$}. These adjacency matrices are randomly initialized with different values and can be dynamically adjusted using gradient backpropagation [12]. Each $\bm{A}^{i}$ can learn one view of graph connections that belongs to a certain basic cognitive process. For each $\Phi_{g}^{i}(\cdot)$, stacking different layers of GCNs can learn different degrees of node cluster similarity [34]. Intuitively, for localized connections, such as electrodes within a brain functional area, a deeper GCN can get a consistent representation among these nodes. However, for global connections that are among different brain functional areas, a shallow GCN can aggregate the information among these areas while not over smooth it. The deeper the $\Phi_{g}^{i}(\cdot)$ is, the taller the feature pyramid level its output has. A linear projection layer, $\textrm{LP}(\cdot)$, is added for each $\Phi_{g}^{i}(\cdot)$ to project the flattened graph representations into a hidden embedding denoted by $\bm{H}_{g}^{i}\in\mathbb{R}^{d_{g}}$. By stacking $\bm{H}_{g}^{i}$ from parallel different-layer GCNs, we can get multi-view pyramid graph embeddings as:
@@ III-C1 Token Mixers for Classification Tasks
Considering the fact that emotion is short-term continuous and long-term varying [17], we propose a short-time aggregation (STA) layer after MSA to learn the long-short-time contextual information. A dropout layer with a scaling factor $\alpha$, $\textrm{dp}(\alpha)$ is applied to control the forgetting rate of the temporal context. $\alpha$ is a hyper-parameter that will scale down the overall dropout rate in STA. To capture short-term consist patterns and smooth the effects of the changes in emotional states, CNN kernels denoted by $\bm{K}_{cnn}$ whose size and step are $(n_{anchor},1)$, $(1,1)$ are utilized to aggregate $n_{anchor}$ temporal neighbors after MSA. Let $\bm{H}_{attn}\in\mathbb{R}^{n_{head}\times seq\times d_{head}}$ denotes one output of MSA. STA can be described as:
@@ V-A Emotion Classification
The observations on the THU-EP and FACED datasets are different. For the THU-EP dataset, EmT-B achieves the best F1 score of 0.724, while Conformer achieves the best accuracy (0.601). On the FACED dataset, EmT-B leads with the second-best accuracy (0.608) and the highest F1 score (0.740). AMDET achieves the second-best F1 score of 0.726, while the best accuracy is achieved by TSception (0.619). As the classes are imbalanced in both the THU-EP and FACED datasets, F1 scores are more important than accuracy. EmT-B achieves the highest F1 scores on both THU-EP and FACED, demonstrating its effectiveness in emotion classification. Different from the observations on the SEED dataset, the baselines using EEG as input achieve better performance than those using features as input. This might be because there are more subjects in THU-EP and FACED, and using EEG directly can provide more information.
For the SEED dataset (62 channels, 15 subjects), there is a positive correlation between the number of transformer layers and performance, with EmT-D (8 layers) performing best due to the enriched spatial information from the higher number of channels. However, for the THU-EP and FACED datasets (32 channels, more subjects), EmT-B (4 layers) achieves better results, likely due to fewer channels and higher subject variability, where deeper models like EmT-D tend to overfit. Overall, EmT-B strikes a better balance between temporal and spatial learning in datasets with fewer channels and greater inter-subject variability.
@@ V-D Effect of EEG Features
Fig. 3 illustrates the impact of different feature types on the SEED dataset for classification tasks and the MAHNOB-HCI dataset for regression tasks. We compared three types of features: PSD, DE, and rPSD, evaluating their effects on accuracy and F1 score for classification tasks, as well as on MRSE, CCC, and PCC for regression tasks. For classification tasks, rPSD outperformed the other two feature types, providing the highest accuracy and F1 score. Specifically, rPSD achieved 5.9% and 11.5% higher accuracy and F1 score, respectively, compared to EmT using DE. Furthermore, rPSD showed a 13.0% and 18.9% improvement in accuracy and F1 score, respectively, over PSD. These results suggest that rPSD is a superior spectral feature for EEG emotion classification tasks. For regression tasks, using rPSD resulted in better MRSE and CCC compared to using DE. However, the difference in PCC between rPSD and DE was minimal, with rPSD achieving a PCC of 0.491 and DE achieving a PCC of 0.490. Notably, when PSD was used as a feature, the model failed to converge, so we excluded these results from further analysis. Based on these findings, rPSD is demonstrated to be a more effective feature than DE, and both rPSD and DE are superior to PSD for EEG emotion recognition tasks.
@@
Figure 4: Effect of the depth (a) and width (b) of GCNs in RMPG on classification performance using SEED. For the depth analysis, 1,1 indicates that both GCN branches in RMPG have 1 layer each. The width is the hidden size of GCN layer.
V-E Effect of The Depth and Width of GCNs in RMPG
We varied the number of GCN layers (depth) and the hidden size of the GCN layers (width) to evaluate their effects on classification performance. EmT-D and the SEED dataset were used for these experiments as they provide better overall performance compared to other configurations. The results are shown in Fig. 4. Since the RMPG module has two GCN branches, we tested four different configurations of GCN layers: [1, 1], [1, 2], [2, 2], and [2, 3]. These numbers represent the number of GCN layers (depth) for each branch in RMPG. As shown in Fig. 4 (a), the [1, 2] configuration yields the best performance. Compared to [1, 1] and [2, 2], the results highlight the effectiveness of the pyramid design. Additionally, increasing the number of GCN layers leads to a significant drop in classification performance, consistent with the over-smoothing issue in deeper GCNs [37]. In contrast, the effects of width are relatively smaller, as shown in Fig. 4 (b). There is a positive correlation between width and performance when increasing the width from 8 to 32. However, the performance declines when the width is increased further, likely due to over-fitting caused by the larger model size.
V-F Effect of The Number of TCT Blocks
We vary the number of TCT blocks from 2 to 8 and monitor their effect on the results of the two tasks. The results are shown in Fig. 5. For classification, shown in Fig. 5 (a), as the number of TCT blocks increases from 2 to 8, there is a noticeable improvement in both ACC and F1 scores. Specifically, accuracy increases from 0.780 to 0.802, while the F1 score shows a more substantial rise from 0.759 to 0.821, with the most significant improvement occurring between 6 and 8 TCT blocks. This indicates that adding more TCT blocks enhances the model’s ability to capture temporal contextual information, thereby improving classification performance. Conversely, for regression tasks, shown in Fig. 5 (b), the number of TCT blocks has little to no effect on the performance metrics. The RMSE remains stable around 0.06, and both the PCC and CCC show minimal variation, hovering around 0.48 and 0.39, respectively. This suggests that while increasing TCT blocks benefits classification by improving the capture of temporal contextual information, it does not significantly impact regression performance.
@@ V-F Effect of The Number of TCT Blocks
V-G Analysis of The Computational Complexity
We conducted an analysis of the computational complexity of various models using floating-point operations (FLOPs) and the number of parameters as evaluation metrics. For the SEED dataset, ACC is used since the classes are balanced, while the F1 score is used for the FACED dataset due to its imbalanced classes. Since THU-EP and FACED share the same data format, we evaluated results only on FACED, which is larger than THU-EP. The results are shown in Fig 7. TSception and Conformer are excluded from the comparison due to their significantly higher FLOPs than the other models. EmT-D achieves the highest ACC on SEED with relatively comparable computational complexity. EmT-B demonstrates higher overall decoding performance across the three datasets with a balance between performance and computational cost.